A copper wire of diameter 1 mm and length 30 m is connected across a battery of 2V.
Calculate the current density in the wire and drift velocity of the electrons. The resistivity
of copper is 1.72 ×10^–8Ωm and n = 8.0 ×10^28 electrons m^–3.
.
1
Expert's answer
2020-02-21T09:59:56-0500

a) Let’s first find the resistance of the copper wire from the formula:


"R = \rho \dfrac{l}{A},"

here, "\rho" is the resistivity of the copper wire, "l" is the length of the wire, "A = \dfrac{\pi d^2}{4}" is the cross-sectional area of the wire and "d" is the diameter of the wire.

Then, we get:


"R = \rho \dfrac{l}{A} = \rho \dfrac{4l}{\pi d^2},""R = 1.72 \cdot 10^{-8} \Omega \cdot m \cdot \dfrac{4 \cdot 30m}{\pi \cdot (1.0 \cdot 10^{-3}m)^2} = 0.65 \Omega."

Then, from the Ohm’s law we can find the current flowing through the copper wire:


"I = \dfrac{V}{R} = \dfrac{2.0 V}{0.65 \Omega} = 3.1A."

Finally, we can find the current density in the wire:


"J = \dfrac{I}{A} = \dfrac{4I}{\pi d^2} = \dfrac{4 \cdot 3.1A}{\pi \cdot (1.0 \cdot 10^{-3}m)^2} = 3.95 \cdot 10^6 \dfrac{A}{m^2}."

b) We can find the drift velocity of the electrons from the formula:


"v = \dfrac{I}{nAq},"

here, "I" is the current flowing through the wire, "n" is the number of free electrons per unit volume of the copper wire, "A" is the cross-sectional area of the wire and "q" is the charge on each electron. Then, we get:


"v = \dfrac{I}{nAq} = \dfrac{4I}{n\pi d^2 q},"

"v = \dfrac{4 \cdot 3.1A}{8.0 \cdot 10^{28} \dfrac{electrons}{m^3} \cdot \pi \cdot (1.0 \cdot 10^{-3}m)^2 \cdot 1.6 \cdot 10^{-19}C} = 3.08 \cdot 10^{-4} \dfrac{m}{s}."

Answer:

a) "J = 3.95 \cdot 10^6 \dfrac{A}{m^2}."

b) "v = 3.08 \cdot 10^{-4} \dfrac{m}{s}."